OpenStudy (anonymous):

If the side length of the cube shown here is changed by a scale factor of 5, what will be the volume of the new cube? A cube with side length of four inches is shown. three hundred twenty cubic inches four hundred cubic inches nine hundred sixty cubic inches eight thousand cubic inches

OpenStudy (anonymous):

\[4^3=64\] \[(5*4)^3=20^3=8000inches^3\]